Let λ be a real number. Suppose that if ABCD is any convex cyclic quadrilateral such that AC = 4, BD = 5, and AB ⊥ CD, then the area of ABCD is at least λ. Then the greatest possible value of λ is m/n, where m and n are positive integers with gcd(m,n) = 1. Compute 100m+n.
1
Expert's answer
2020-12-16T20:26:45-0500

The area of a quadrilateral with sides a, b, c, d:

"S=\frac{1}{2}(ac+bd)sin\theta" , where "\theta" is an angle between diagonals

By Ptolemy's Theorem:

"ac+bd=ef" , where e, f are diagonals

In our case:

"e=AC, f=BD"

Then:

"S=\frac{4\cdot5}{2}sin\theta=10sin\theta"

"S_{max}=10sin90\degree=10=m/n"

"m=10, n=1" since "gcd(m,n) = 1"

So:

"100m+n=100\cdot10+1=1001"